OpenStudy (anonymous):

The cost of four scarves and six hats is $52.00. The cost of two hats is $1.00 more than the cost of one scarf. What is the cost of one scarf?

hero (hero):

4s+6h = 52 2h = s+1

hero (hero):

s = 13-1.5h s = 2h - 1 s = s 13 - 1.5h = 2h - 1 14 = .5h 7 = h s = 13